# Friendship-paradox paradox: Do most people’s friends really have more friends than they do?

### v4

@@ II.2 Classical alter-based and ego-based FP inequalities
Those are the original formulations presented in Ref. [1]. Recent works [2, 3] have shown that the difference between the left-hand sides of Eqs. (5) and (6) is determined by the covariance between $k_{i}$ and $k_{nn}(i)$ over nodes [3], i.e.,
$\langle k_{\text{friend}}\rangle_{\mathrm{n}}-\langle k_{nn}\rangle_{\mathrm{n}}=\frac{1}{\langle k\rangle_{\mathrm{n}}}\operatorname{Cov_{\mathrm{n}}}(k,k_{nn})\,$, (7)
(verified in Table 1 for the example networks analyzed in this work) or equivalently (the equivalence is proved in Ref. [3]) as presented in Ref. [2],
$\langle k_{nn}\rangle_{\mathrm{n}}-\langle k_{\text{friend}}\rangle_{\mathrm{n}}=\rho\sqrt{\left(\frac{\kappa_{1}\kappa_{3}-\kappa_{2}^{2}}{\kappa_{1}}\right)\left(\kappa_{-1}-\kappa_{1}^{-1}\right)}\,$, (8)
where $\rho=\mathrm{Corr_{\mathrm{e}}}\left(D_{D},D_{O}^{-1}\right)$, which is the Pearson correlation coefficient between the degree $D_{D}$ of the origin node (“the origin degree” in Ref. [2]) and the reciprocal of the degree $D_{O}$ of the destination node (“the inverse destination degree” in Ref. [2]) and $\kappa_{m}=\left(\sum_{i=1}^{N}k_{i}^{m}\right)/N$ is the $m$th moment of degrees.
@@ II.4 Local median-based majority fraction
In the sections that follow, these quantities are evaluated in a simple toy network, in the ZKC network, and in the AFB network, and their joint behavior is interpreted in terms of their local organization of joint distributions of $h_{i}$ and $k_{nn}$.
@@ IV Empirical Examples: Karate Club and American Football Networks
To illustrate the four measures discussed in this paper, consider two simple (and perhaps overused, but not in this context) empirical data: Zachary’s karate club (ZKC) network [9] and the National Collegiate Athletic Association Division I American football (AFB) network compiled and published in Refs. [10, 11]. For each network, the two classical FP quantities $\langle k_{\mathrm{friend}}\rangle_{\mathrm{n}}$ and $\langle k_{nn}\rangle_{\mathrm{n}}$ are computed, along with the two median-based fractions defined earlier:
$\phi_{\mathrm{global}}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{\{k_{i}<k_{nn}(i)\}}\,,\qquad\phi_{\mathrm{local}}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{\{h_{i}<1/2\}}\,.$ (23)
Table 1 summarizes the numerical values of the four quantities for all three networks.

### v5

@@ II.2 Classical alter-based and ego-based FP inequalities
Those are the original formulations presented in Ref. [1]. For completeness, we briefly outline how the relation between the two averages arises.
The alter-based mean degree can be written as
$\langle k_{\mathrm{friend}}\rangle_{\mathrm{n}}=\frac{\langle k^{2}\rangle_{\mathrm{n}}}{\langle k\rangle_{\mathrm{n}}}\,$, (7)
which follows from the fact that sampling nodes via edges biases the degree distribution by a factor proportional to $k$. On the other hand, the ego-based mean is obtained by averaging the mean neighbor degree
$k_{nn}(i)=\frac{1}{k_{i}}\sum_{j}a_{ij}k_{j}\,$, (8)
over all nodes,
$\langle k_{nn}\rangle_{\mathrm{n}}=\frac{1}{N}\sum_{i}k_{nn}(i)\,.$ (9)
A key combinatorial identity connects these quantities. Multiplying $k_{nn}(i)$ by $k_{i}$ and summing over all nodes yields
$\sum_{i}k_{i}k_{nn}(i)=\sum_{i,j}a_{ij}k_{j}=\sum_{i}k_{i}^{2}\,$, (10)
where the second equality follows from the symmetry $a_{ij}=a_{ji}$. Dividing by $N$ gives
$\langle kk_{nn}\rangle_{\mathrm{n}}=\langle k^{2}\rangle_{\mathrm{n}}\,.$ (11)
Substituting this relation into the definition of $\langle k_{\mathrm{friend}}\rangle_{\mathrm{n}}$ immediately yields
$\langle k_{\mathrm{friend}}\rangle_{\mathrm{n}}=\frac{\langle kk_{nn}\rangle_{\mathrm{n}}}{\langle k\rangle_{\mathrm{n}}}\,.$ (12)
Subtracting $\langle k_{nn}\rangle_{\mathrm{n}}$ from both sides then gives
$\langle k_{\mathrm{friend}}\rangle_{\mathrm{n}}-\langle k_{nn}\rangle_{\mathrm{n}}=\frac{1}{\langle k\rangle_{\mathrm{n}}}\operatorname{Cov}_{\mathrm{n}}(k,k_{nn})\,$, (13)
where the node-based covariance $\operatorname{Cov_{\mathrm{n}}}(k,k_{nn})=\langle k\,k_{nn}\rangle_{\mathrm{n}}-\langle k\rangle_{\mathrm{n}}\langle k_{nn}\rangle_{\mathrm{n}}$ indicates the covariance of the degree of a node and the average degree of its neighbors. This identity shows that the difference between the alter-based and ego-based friendship paradoxes is governed entirely by the covariance between a node’s degree and the average degree of its neighbors.
Equation (13) can be rewritten in an equivalent moment-based form originally derived in Ref. [2] and shown to be analytically identical in Ref. [3],
$\langle k_{nn}\rangle_{\mathrm{n}}-\langle k_{\mathrm{friend}}\rangle_{\mathrm{n}}=\rho\sqrt{\left(\frac{\kappa_{1}\kappa_{3}-\kappa_{2}^{2}}{\kappa_{1}}\right)\left(\kappa_{-1}-\kappa_{1}^{-1}\right)}\,$, (14)
where $\rho=\mathrm{Corr}_{\mathrm{e}}(D_{D},D_{O}^{-1})$ is the Pearson correlation coefficient, measured over oriented edges, between the degree $D_{D}$ of the destination node and the inverse degree $D_{O}^{-1}$ of the origin node. Here $\kappa_{m}=N^{-1}\sum_{i=1}^{N}k_{i}^{m}$ denotes the $m$th moment of the degree distribution. This representation makes explicit how the difference between the two friendship paradox averages depends not only on degree heterogeneity (through the degree moments $\kappa_{m}$) but also on edge-level correlations encoded in $\rho$. While Eq. (13) emphasizes a node-level covariance between $k_{i}$ and $k_{nn}(i)$, Eq. (14) expresses the same quantity in terms of edge-sampled statistics, thereby highlighting the role of degree–degree correlations along edges. The equivalence between Eqs. (13) and (14) thus clarifies that both expressions capture the same underlying structural mechanism, viewed from complementary node-based and edge-based perspectives.
@@ II.4 Local median-based majority fraction
In the sections that follow, these quantities are evaluated in a simple toy network, in the Zachary’s karate club (ZKC) network [9], and in the National Collegiate Athletic Association Division I American football (AFB) network [10, 11], and their joint behavior is interpreted in terms of their local organization of joint distributions of $h_{i}$ and $k_{nn}$.
@@ IV Empirical Examples: Karate Club and American Football Networks
To illustrate the four measures discussed in this paper, consider two simple (and perhaps overused, but not in this context) empirical data: the ZKC network [9] and the AFB network compiled and published in Refs. [10, 11]. For each network, the two classical FP quantities $\langle k_{\mathrm{friend}}\rangle_{\mathrm{n}}$ and $\langle k_{nn}\rangle_{\mathrm{n}}$ are computed, along with the two median-based fractions, $\phi_{\mathrm{global}}$ and $\phi_{\mathrm{local}}$, defined earlier in Eqs. (15) and (21). Table 1 summarizes the numerical values of the four quantities for all three networks.